OpenStudy (anonymous):

Simplify: √28 over 25 @greenlegodude57

OpenStudy (anonymous):

A. \[\frac{4√7 }{ 25 } \]

OpenStudy (anonymous):

\[ \sqrt{3}\]

OpenStudy (anonymous):

C. \[\frac{ 4 \sqrt{7} }{ 25 }\]

OpenStudy (anonymous):

D. \[\frac{ 2\sqrt{7} }{ 5 }\]

OpenStudy (anonymous):

@karategirl2002 @greenlegodude57

OpenStudy (anonymous):

i think C

OpenStudy (karategirl2002):

1.0583000524

OpenStudy (anonymous):

?

OpenStudy (karategirl2002):

a

OpenStudy (anonymous):

oh A has the denominator of 5 not 25

OpenStudy (anonymous):

Answer is D.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Oh, sorry. I thought D had a denominator of 25. If it had a denominator of 25, it would be right. But sorry.

OpenStudy (anonymous):

so C is right and its fine :D

OpenStudy (anonymous):

A is right too

OpenStudy (anonymous):

Just saw that you said A had a denominator of 5

OpenStudy (anonymous):

C is wrong

OpenStudy (anonymous):

A has a denominator of 5 i typed it wrong

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

\[\frac{ \sqrt28 }{ 25 } = 0.211660104885\] \[\frac{ 4\sqrt7 }{ 5 } = 0.211660104885\]

OpenStudy (anonymous):

See? They're the same

OpenStudy (anonymous):

oh okay thnks :{)

OpenStudy (anonymous):

\[\frac{ \sqrt{28} }{ 25 }=\frac{ 2\sqrt{7} }{ 25 }\]

OpenStudy (anonymous):

@GodGirl

OpenStudy (anonymous):

@SiennaSurfer its A

OpenStudy (anonymous):

@SiennaSurfer That is correct, but it is NOT one of the choices.

OpenStudy (anonymous):

i didnt see multiple choise. oh well, still doesnt change my answer

OpenStudy (anonymous):

i posted them in the comment section but okay...

OpenStudy (anonymous):

I know, I thought D was \[\frac{ 2\sqrt7 }{ 25 }\] but instead it was \[\frac{ 2\sqrt7 }{ 5 }\]

OpenStudy (anonymous):

thats not right

OpenStudy (anonymous):

omg im confused now

OpenStudy (anonymous):

@greenlegodude57

OpenStudy (anonymous):

The answer is A, the answer @SiennaSurfer put was CORRECT but it was NOT in the multiple choices. Choice A is also correct and IS IN the multiple choices.

OpenStudy (anonymous):

@GodGirl , did u accidentally put one answer twice, a and c r the same

OpenStudy (anonymous):

a is not correct

OpenStudy (anonymous):

A is correct, she said she made a mistake and the bottom number is 5, not 25.

OpenStudy (anonymous):

ik i already have cleard that up. instead of 25 on A its suppose 2 be 5

OpenStudy (anonymous):

@SiennaSurfer

OpenStudy (anonymous):

nvmd. can u tell us if u got it right?